A chunk of plastic is weighed on a scale and has a mass of 53.87 g. Next, the chunk is submerged in a graduated cylinder with an initial volume of water of 100.0 mL. After the plastic is added, the new volume in the cylinder is 113.5 mL. Find the density of the plastic in g/cm3.


1
Expert's answer
2021-04-19T05:14:24-0400

"Density=\frac {mass}{ volume}"

Mass = 53.87g

Volume = 113.5 - 100= 13.5ml

"Density= \frac {53.87}{ 13.5} = 3.99g/cm^3"

Ans= 3.99g/cm³